April 22, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, April 22, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Connection Using Bluejeans

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 630804895.
  2. To join via a Web Browser, go to the page [1] https://bluejeans.com/630804895.
  3. To join via phone, use one of the following numbers and the Conference ID: 630804895
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
  5. More information on connecting to bluejeans is available.


  1. Announcements
  2. Collaboration Meeting
  3. Subdetector Reports
    • Calibration Progress
    • Resolutions / Efficiencies
  4. Simulations
  5. Spring Calibration Plans
  6. Normalization


Attending: Sean, Simon, Adesh, Lubomir, Justin, Paul, Eugene, Will M., Beni, Alex B., Eric, Elton, Mike S., Luke, Kei (JLab); Curtis (CMU)


Adesh reported that not much progressed had been made. The major current puzzle is that when looking at showers due to electrons, although the overall cluster energy has a resolution ~10%, the individual block energies have resolutions of ~30%. They have made progress in determining the efficiencies of showers due to electrons, but are held up with this problem. Will M. has agreed to look at the pi0 minimization procedure, based on his experience with using it for the BCAL.


Elton reported that the efficiencies depend on many variables, including how different types of hits are treated and which layer is being looked at. These issues are a major current focus. Another recent major effort was developing better running conditions. These include variable fan speeds on the crates (h/t Mark Dalton), running at a lower temperature (10 degrees down from 18 degrees), and a higher over-voltage (1.4V versus 1.0V). Will M. is looking for eta -> gamma gamma decays, since none have been seen yet so far in the fall data. He is looking at the data and simulations to determine the prospects for measuring this. It was also suggested that the characterize the pi0 reconstruction performance over a wider energy range.


Naomi has continued looking into the gas mix, and has developed a plugin for monitoring the drift times. David L. and Sean have been looking into running it as an online monitoring process. Mike S. has been looking into using the Millipede program as a complementary alignment strategy. This is a standard program used by many other experiments for alignment, by building and inverting large sparse matrices. For comparison, CMS did the alignment of their silicon tracker muon system, which had ~4000 parameters, using 4M cosmic rays. Over a weekend, with a good DAQ running (rate ~1.8 kHz, live time ~99 %), we recorded ~100M triggers, O(1-5%) of which are cosmic rays.

More data with the HV power supply filters are needed to characterize the FDC, since the noise levels are now back down to normal. Beni said that he had taken good cosmic data lately. Eugene suggested running this procedure on simulated data to make sure it works.


Lubomir has given a full set of alignments to Simon, who is checking them out today. These are only first order corrections - there is some residual sinusoidal dependence (probably due to signal propagation) that remains to be understood. The reconstruction is still only using cathodes, but now the wire TDC times are better understood, so that they can be integrated.


Beni recently made per-channel time offset corrections, and Simon checked them into the CCDB. These are the first offsets that have been added for the TOF. See Tuesday's TOF meeting for the details of the calibration, but note that it was only done for double-ended paddles of normal width (i.e. not the single-ended or narrow paddles around the beam line), and greatly enabled by knowing the correct CAEN TDC time binning.

Simon finds the combined TOF/ST time resolution to have decreased from ~1ns to 700 ps. Beni also obtained effective velocity values that Simon is looking into. Sean is looking at the track/TOF matching in some detail to make sure that these new constants line up with other detectors. Energy scale and updated time walk corrections are also on the list.

Start Counter

Eric is looking at several improvements to the start counter calibrations. He is looking at an improved version of the time walk corrections based on the pulse height instead of the pulse integral, based on Elton's suggestions. There was some discussion of this topic, and the main problems with using the pulse integral is dealing with the effect of reflections in the scintillator and overflows for larger pulses. The time walk corrections depend on the leading edge of the pulse, so the pulse height is more relevant. It was suggested that Eric look at pulses characteristics as a function of position on the ST and to use CDC dE/dx information to study pion versus proton tracks.

Eric is also looking at propogation time corrections, and trying to understand why the results obtained using Fall data differ from the bench measurements at FIU. These numbers are close to the default value in the PID library, so we should be good to wait for the larger statistics needed to refine these calculations.

N.B. Simon updated the energy scale factors for the start counter, FDC, and CDC.


Nathan continues to work on the PS-tagger cross calibration and efficiency determination.

For the PS: Some software changes were needed to support lower energy running: photon beam endpoint energies are now taken from the CCDB, the rough energy range for each counter was refined, and the monitoring plots were updated. He is seeing a time resolution of ~00 for the PSC.


Alex finished his gain studies and is waiting for beam data (with PS) to test them.

RF Timing

Paul M. described the RF timing signal that was recently hooked into TDCs on various crates, and his recent efforts in writing a plugin to monitor and calibrate these signals. He walked us through various plots that he used to calibrate the signal, looking at the time differences between the multiple RF signals read out in each event. One important finding was that the TOF time bin size was not 25 ps, but actually about 7% smaller. The detailed information was unearthed from the CAEN TDC manual. So far, the time resolutions range from ~70 ps (FDC crate) to ~25 ps (TOF crate). The TAGH and PS crates had no data since they were not being rad out in the recent cosmic runs that he was using to calibrate with.

Elton pointed out that the prescaling factors used to read out these pulses are actually fixed, but a fix is in the work to allow them to be set in firmware.


The cable swap in the translation table that was fixed a couple of months ago never made it into the CCDB, and this was reflected in the latest offline monitoring results. Paul M. tracked this down and fixed it.

He made a couple other notes: TOF matching to individual paddles in the two planes was added; Start counter matching is good at large angles but not in the forward direction, which is unsurprising since the uncertainties in the tracks are larger, and the paddles get smaller.

Sean is adding calibration and efficiency information to the monitoring DB so that we can track it.


Simon has made some improvements and fixes to the geometry. Sean has run some checks and pronounces the MC fit to run. He is waiting to push the button and flood the queue with jobs to talk to Mark Ito and finalize the conditions.

We reviewed the current status of the smearing parameters in mcsmear, to prepare for large scale simulations in the summer. Most of these can be updated once higher quality data or the outstanding problems described earlier can be solved. For the BCAL, which has the most complicated smearing, Tegan is working to replace the detailed calculations that depend on the pulse shape with a simpler parameterization. The stability studies also help in determining the noise level to use for MC. Generally it was pointed out that addition of noise hits is turned off by default - the issue overlaps somewhat with the idea of overlaying EM background events that Sean is looking into (see last week's offline software meeting).